# font=, innertopmargin=2pt, innerbottommargin=2pt, innerleftmargin=8pt, innerrightmargin=8pt, frametitleaboveskip=2pt, frametitlebelowskip=1pt, frametitlealignment=, middlelinecolor=blue

### v2

@@ 1 Introduction
While advanced data-driven learning-based methods offer various solutions to generate synthetic individual trajectories [13, 39, 9, 6], the generated data only imitates real-world data from the data distribution perspective rather than semantics, rendering them less effective in simulating or interpreting activities in novel or unforeseen scenarios with a significantly different distribution (e.g., a pandemic). Thus, in this study, to explore a more intelligent and effective activity generation, we propose to establish a trajectory generation framework by exploiting the emerging intelligence of LLM agents, as illustrated in Figure 1. LLMs present two significant advantages over previous models when applied to activity trajectory generation:
  • Semantic Interpretability. Unlike previous models, which have predominantly depended on structured data (e.g., GPS coordinates-based trajectory data) for both calibration and simulation [15, 26], LLMs exhibit proficiency in interpreting semantic data (e.g., activity trajectory data). This advantage significantly broadens the scope for incorporating a diverse array of data sources into generation processes, thereby enhancing the models’ ability to understand and interact with complex, real-world scenarios in a more nuanced and effective manner.
@@
To this end, our study employs LLM agents to infer activity patterns and motivation for personal activity generation tasks. While previous researches advocate habitual activity patterns and motivations as two critical elements for activity generation [15, 38], our proposed framework introduces a more interpretable and effective solution. By leveraging the capabilities of LLMs to process semantically rich datasets (e.g., personal check-in data), we enable a nuanced and interpretable simulation of personal mobility. Our methodology revolves around two phases: (1) activity pattern identification and (2) motivation-driven activity generation. In Phase 1, we leverage the semantic awareness of LLM agents to extract and identify self-consistent, personalized habitual activity patterns from historical data. In Phase 2, we develop two interpretable, retrieval-augmented strategies that utilize the patterns identified in Phase 1. These strategies guide LLM agents to infer underlying daily motivations, such as evolving interests or situational needs. Finally, we instruct LLM agents to play urban residents according to the obtained patterns and motivations and generate their daily activities.
@@ 2.1 Personal Mobility Generation
Activity trajectory generation offers a valuable perspective for understanding personal mobility. Based on vast call detailed records, Jiang et al. built a mechanistic modeling framework to generate individual activities in high spatial-temporal resolutions. Pappalardo and Simini employed Markov modeling to estimate the probability of individuals visiting specific locations. Besides, deep learning has become a robust tool for modeling the complex dynamics of traffic  [18, 13, 39, 9, 6, 20]. The primary challenge involves overcoming data-related obstacles such as randomness, sparsity, and irregular patterns [8, 38, 37, 19]. For example, Feng et al. proposed attentional recurrent networks to handle personal preference and transition regularities. Yuan et al. leveraged deep learning combined with neural differential equations to address the challenges of randomness and sparsity inherent in irregularly sampled activities for activity trajectory generation. Recently,  Zhu et al. proposed to utilize a diffusion model to generate GPS trajectories.
@@ 2.2 LLM Agents in Social Science
Exploring how to treat LLMs as autonomous agents in specific scenarios lead to diverse and promising applications in social science [30, 34, 33, 10]. For instance, Park et al. established an LLM-powered agent framework to simulate human behavior in an interactive scenario, demonstrating the potential of LLMs to model complex social interactions and decision-making processes. Moreover, the application of LLM agents in economic research has been explored, providing new insights into financial markets and economies [11, 17]. Extending beyond the realm of social sciences, Mao et al. adeptly utilized LLMs to generate driving trajectories in motion planning tasks. In the field of natural sciences, Williams et al. integrated LLMs with epidemic models to simulate the spread of diseases. These varied applications highlight the versatility and potential of LLMs to understand and model various real-world dynamics.
@@ 3 Methodology
By modeling individuals within an urban environment as LLM agents, we present LLMob, an LLM Agent Framework for Personal Mobility Generation, as illustrated in Figure 2. LLMob is based on the assumption that an individual’s activities are primarily influenced by two principal factors: habitual activity patterns and current motivations. Habitual activity patterns, representing typical movement behaviors and preferences that indicate regular travel and location choices, are recognized as crucial information for inferring daily activities [7, 29]. On the other hand, motivations relate to dynamic and situational elements that sway an individual’s choices at any particular moment, such as immediate needs or external circumstances during a specific period. This consideration is vital for capturing and forecasting short-term shifts in mobility patterns [1, 38]. Moreover, by formulating prompts that assume specific events of concern, this framework allows us to observe the LLM agent’s responses in a variety of situations.
To construct a pipeline for activity trajectory generation, we design an LLM agent with action, memory, and planning [31, 30]. Action specifies how an agent interacts with the environment and makes decisions. In LLMob, the environment contains the information collected from real-world data, and the agent act by generating trajectories. Memory includes past actions that need to be prompted to the LLM to invoke the next action. In LLMob, memory refers to the patterns and motivations output by the agent. Planning formulates or refines a plan over past actions to handle complex tasks, with additional information optionally incorporated as feedback. In LLMob, we use planning to identify patterns and motivations, thereby handling the complex task of trajectory generation. Plan formulation, selection, reflection, and refinement [14] are employed in succession, and the agent keeps updating the action plan based on its observation [36]: The agent first formulates a set of activity plans by extracting candidate patterns from historical trajectories in the database. Then, the agent performs self-reflection through a self-consistency evaluation to pick the best pattern from the candidate patterns. With historical trajectories further retrieved from the database, the agent refines the identified pattern to a summarized motivation of daily activity, which is then jointly used with the identified pattern for trajectory generation. Besides the above agentic components, we also specify personas to the agent to simulate the diversity of real-world individuals [28].
@@ 3.1.1 Pattern Extraction from Semantics and Historical Data
This step derives activity patterns based on activity trajectory data (e.g. individual check-in data). As illustrated in the left panel of Figure 2, this scheme consists of the following aspects: For each person, we start by specifying a candidate personas to the LLM agent, providing the inspiring foundation for subsequent activity pattern generation. This approach also encourages the diversity of the generated activity patterns, as each candidate persona acts as a unique prior for the generation process (e.g., the significance of user clustering from activity trajectory data in producing meaningful distinctions has been demonstrated [23]). Meanwhile, we perform data preprocessing to extract key information from the extensive historical data. This involves identifying usual commuting distances, pinpointing typical start and end times and locations of daily trips, and concluding the most frequently visited locations of the person. It is important to note that these pieces of information are widely recognized as critical features in mobility analysis [15]. Following this, both semantic elements with historical data are combined in the prompts, requiring the LLM agent to summarize the activity patterns for this person. By doing this, we set up a streamline to effectively bridge the gap between semantic persona characteristics and concrete historical activity trajectory data, which allows for a more personalized and interpretable representation of individual activities at one day. For each person, a set of $C$ ($C=10$) candidate patterns, denoted as $\mathcal{CP}$, are generated according to the historical data and $C$ candidate personas, respectively.
@@ 3.1.2 Pattern Evaluation with Self-Consistency
This step involves assessing the consistency of the candidate patterns to identify the most plausible one. We implement a scoring mechanism to gauge the alignment of candidate patterns with historical data. To achieve this objective, we define a scoring function for each candidate pattern $cp$ in the set $\mathcal{CP}$. This function evaluates $cp$ against two distinct sets of activity trajectories: the specific activity trajectories $\mathcal{T}_{i}$ of a targeted resident $i$ and the sampled activity trajectories from other residents $\mathcal{T}_{\sim i}$:
$score_{cp}=\sum_{t\in\mathcal{T}_{i}}r_{t}-\sum_{t^{\prime}\in\mathcal{T}_{\sim i}}r_{t^{\prime}}$ (1)
where we design an evaluation prompt to ask the LLM to generate rating scores $r_{t}$ and $r_{t^{\prime}}$. Specifically, the LLM agent is prompted to assess the degree of preference for a given trajectory based on the candidate pattern. Ideally, the LLM agent should assign a higher $r_{t}$ for data from the targeted resident and a lower $r_{t^{\prime}}$ for data from other residents. This scheme requires self-consistency in the pattern: the activity pattern derived from the activity trajectory data should be consistent with the data during the evaluation. We provide the pseudo-code of the algorithm for Phase 1 of LLMob in Appendix A.
@@ 3.2 Motivation-Driven Activity Generation
In Phase 2 of LLMob, we focus on the retrieval of motivation and the integration of motivation and activity patterns for individual activity trajectory generation. Retrieval-augmented generation has been identified as a crucial factor in boosting the performance of LLM [35]. This enhancement provides additional information that aids LLM in more effectively responding to queries. While previous studies on activity generation mainly overlook the critical factors of macro temporal information (e.g., date) or specific scenarios (e.g., harsh weather) [37], we propose a more sophisticated activity generation which accounts for various conditions by taking advantage of the human-like intelligence of LLM. For instance, the activity trajectory at date $d$ can be inferred given the motivation of this date and the habitual activity pattern as:
@@
This generation scheme instructs the LLM agent to simulate a designated individual according to a given activity pattern, and then meticulously generate an activity trajectory in accordance with the daily motivation. To obtain insightful and reliable motivations, two retrieval schemes are proposed.
@@ 3.2.1 Evolving-based Motivation Retrieval
This scheme is related to the intuitive principle that an individual’s motivation on any given day is influenced by her interests and priorities in preceding days [27]. Guided by this understanding, our approach harnesses the intelligence of the LLM agent to understand the behavior of daily activities and the underlying motivations. As illustrated in Figure 3, for a specific date $d$ for which we aim to generate the activity trajectory, we consider the activities of the past $k$ days ($k=\min(7,l)$, where $l$ is the maximum value such that the trajectory for date $d-l$ can be found in the database), and prompt the LLM agent to act as an urban resident based on the pattern identified in Section 3.1 and summarize $k$ motivations behind these activities. Leveraging these summarized motivations, the LLM agent is further prompted to infer potential motivation for date $d$.
@@ 3.2.2 Learning-based Motivation Retrieval
$sim_{d_{a},d_{b}}=\sum_{t=1}^{N_{d}}\mathbf{1}_{(\mathcal{T}_{d_{a}}(t)=\mathcal{T}_{d_{b}}(t))}\text{ if }|\mathcal{T}_{d_{a}}|>t\text{ and }|\mathcal{T}_{d_{b}}|>t$ (3)
@@
    $\mathcal{L}(\theta)=\sum_{n=1}^{N}-\log\left(\frac{e^{\text{logits}_{i}}}{\sum^{k+1}_{j=1}e^{\text{logits}_{j}}}\right)_{n}$ (5)
@@
Upon training a similarity score approximation model, it can be applied to access the similarity between any given query date and historical dates. This enables us to retrieve the most similar historical data, which is prompted to the LLM agent to generate a summary of the motivations prevalent at that time. By doing so, we can extrapolate a motivation relevant to the query date, providing a basis for the LLM agent to generate a new activity trajectory.
@@ 4.1 Experimental Setup
Dataset. We investigate and validate LLMob over a personal activity trajectory dataset from Tokyo. This dataset was obtained through Twitter and Foursquare APIs and covers the data from January 2019 to December 2022. The time frame of this dataset is insightful as it captures typical daily life prior to the COVID-19 pandemic (i.e., normal period) and subsequent alterations during the pandemic (i.e., abnormal period). To facilitate a cost-efficient and detailed analysis for different periods, we randomly choose 100 users to model their individual activity trajectory at a 10-minute interval according to the number of available trajectories (samples given in Appendix C.2). We utilize the category classification in Foursquare to determine the activity category for each location. We use 10 candidate personas (Appendix C.4) as a prior for subsequent pattern generation, which captures a diverse range of activity patterns within the data of this study. For the application to other datasets, this style of candidate patterns can be easily initialized using an LLM.
Metrics. The following features related to personal activity are used as metrics: (1) Step distance (SD)[37]: This metric evaluates the spatial pattern of an individual’s activities by measuring the distance between two consecutive locations in a trajectory. (2) Step interval (SI)[37]: This metric evaluates the temporal pattern of an individual’s activities by measuring the time interval between two successive locations on an individual’s trajectory. (3) Daily activity routine distribution (DARD): This metric captures and examines the distribution of activities for each trajectory, focusing on the patterns of individual activities characterized by activity type and timing (e.g., activity type, time). It provides insight into how activities are distributed over space and time and reflects semantic information such as habitual behavior. (4) Spatial-temporal visits distribution (STVD): This metric provides a granular perspective on the generated activities by assessing the spatial-temporal distribution of visited locations within each trajectory, including geographical coordinates and timestamps. It enables a detailed analysis of where and when activities occur.
After extracting the above features from both the generated and real trajectory data, Jensen-Shannon divergence (JSD) is employed to quantify the discrepancy between them. Lower JSD is preferred.

### v3

@@ 1 Introduction
While advanced data-driven learning-based methods offer various solutions to generate synthetic individual trajectories [13, 39, 9, 6, 16, 38], the generated data only imitates real-world data from the data distribution perspective rather than semantics, rendering them less effective in simulating or interpreting activities in novel or unforeseen scenarios with a significantly different distribution (e.g., a pandemic). Thus, in this study, to explore a more intelligent and effective activity generation, we propose to establish a trajectory generation framework by exploiting the emerging intelligence of LLM agents, as illustrated in Figure 1. LLMs present two significant advantages over previous models when applied to activity trajectory generation:
  • Semantic Interpretability. Unlike previous models, which have predominantly depended on structured data (e.g., GPS coordinates-based trajectory data) for both calibration and simulation [17, 27, 47], LLMs exhibit proficiency in interpreting semantic data (e.g., activity trajectory data). This advantage significantly broadens the scope for incorporating a diverse array of data sources into generation processes, thereby enhancing the models’ ability to understand and interact with complex, real-world scenarios in a more nuanced and effective manner.
@@
To this end, our study employs LLM agents to infer activity patterns and motivation for personal activity generation tasks. While previous researches advocate habitual activity patterns and motivations as two critical elements for activity generation [17, 43], our proposed framework introduces a more interpretable and effective solution. By leveraging the capabilities of LLMs to process semantically rich datasets (e.g., personal check-in data), we enable a nuanced and interpretable simulation of personal mobility. Our methodology revolves around two phases: (1) activity pattern identification and (2) motivation-driven activity generation. In Phase 1, we leverage the semantic awareness of LLM agents to extract and identify self-consistent, personalized habitual activity patterns from historical data. In Phase 2, we develop two interpretable retrieval-augmented strategies that utilize the patterns identified in Phase 1. These strategies guide LLM agents to infer underlying daily motivations, such as evolving interests or situational needs. Finally, we instruct LLM agents to act as urban residents according to the obtained patterns and motivations. In this way, we generate their daily activities in a specific reasoning logic.
@@ 2.1 Personal Mobility Generation
Activity trajectory generation offers a valuable perspective for understanding personal mobility. Based on vast call detailed records,  Jiang et al. built a mechanistic modeling framework to generate individual activities in high spatial-temporal resolutions. Pappalardo and Simini employed Markov modeling to estimate the probability of individuals visiting specific locations. Besides, deep learning has become a robust tool for modeling the complex dynamics of traffic  [15, 13, 44, 9, 21]. The primary challenge involves overcoming data-related obstacles such as randomness, sparsity, and irregular patterns [8, 43, 42, 20]. For example, Feng et al. proposed attentional recurrent networks to handle personal preference and transition regularities. Yuan et al. leveraged deep learning combined with neural differential equations to address the challenges of randomness and sparsity inherent in irregularly sampled activities for activity trajectory generation. Recently,  Zhu et al. proposed to utilize a diffusion model to generate GPS trajectories.
@@ 2.2 LLM Agents in Social Science
Exploring how to treat LLMs as autonomous agents in specific scenarios leads to diverse and promising applications in social science [32, 36, 35, 10]. For instance, Park et al. established an LLM agent framework to simulate human behavior in an interactive scenario, demonstrating the potential of LLMs to model complex social interactions and decision-making processes. Moreover, the application of LLM agents in economic research has been explored, providing new insights into financial markets and economies [11, 19]. Extending beyond the realm of social sciences, Mao et al. adeptly utilized LLMs to generate driving trajectories in motion planning tasks. In the field of natural sciences,  Williams et al. integrated LLMs with epidemic models to simulate the spread of diseases. These varied applications highlight the versatility and potential of LLMs to understand and model various real-world dynamics.
@@ 3 Methodology
By modeling individuals within an urban environment as LLM agents, we present LLMob, an LLM Agent Framework for Personal Mobility Generation, as illustrated in Figure 2. LLMob is based on the assumption that an individual’s activities are primarily influenced by two principal factors: habitual activity patterns and current motivations. Habitual activity patterns, representing typical movement behaviors and preferences that indicate regular travel and location choices, are recognized as crucial information for inferring daily activities [31, 7, 30]. On the other hand, motivations relate to dynamic and situational elements that sway an individual’s choices at any particular moment, such as immediate needs or external circumstances during a specific period. This consideration is vital for capturing and forecasting short-term shifts in mobility patterns [1, 43]. Moreover, by formulating prompts that assume specific events of concern, this framework allows us to observe the LLM agent’s responses in a variety of situations.
To construct a pipeline for activity trajectory generation, we design an LLM agent with action, memory, and planning [33, 32]. Action specifies how an agent interacts with the environment and makes decisions. In LLMob, the environment contains the information collected from real-world data, and the agent acts by generating trajectories. Memory includes past actions that need to be prompted to the LLM to invoke the next action. In LLMob, memory refers to the patterns and motivations output by the agent. Planning formulates or refines a plan over past actions to handle complex tasks, with additional information optionally incorporated as feedback. In LLMob, we use planning to identify patterns and motivations, thereby handling the complex task of trajectory generation. Plan formulation, selection, reflection, and refinement [14] are employed in succession, and the agent keeps updating the action plan based on its observation [41]: The agent first formulates a set of activity plans by extracting candidate patterns from historical trajectories in the database. The agent then performs self-reflection through a self-consistency evaluation to pick the best pattern from the candidate patterns. With historical trajectories further retrieved from the database, the agent refines the identified pattern to a summarized motivation of daily activity, which is then jointly used with the identified pattern for trajectory generation. In addition to the above agentic components, we also suggest the personas of the agent, which can facilitate the LLM to simulate the diversity of real-world individuals [29].
@@ 3.1.1 Pattern Extraction from Semantics and Historical Data
This step derives activity patterns based on activity trajectory data (e.g., individual check-in data). As illustrated in the left panel of Figure 2, this scheme consists of the following aspects: For each person, we start by specifying a candidate personas to the LLM agent, providing the inspiring foundation for subsequent activity pattern generation. This approach also encourages the diversity of the generated activity patterns, as each candidate persona acts as a unique prior for the generation process (e.g., the significance of user clustering from activity trajectory data in producing meaningful distinctions has been demonstrated [24]). Meanwhile, we perform data preprocessing to extract key information from the extensive historical data. This involves identifying usual commuting distances, pinpointing typical start and end times and locations of daily trips, and concluding the most frequently visited locations of the person. It is important to note that these pieces of information are widely recognized as critical features in mobility analysis [17]. After the preprocessing procedure, both semantic elements with historical data are combined in the prompts, requiring the LLM agent to summarize the activity patterns for this person. By doing this, we set up a streamline to effectively bridge the gap between semantic persona characteristics and concrete historical activity trajectory data, which allows for a more personalized and interpretable representation of individual activities in one day. Moreover, we propose adding candidate personas to the prompt during candidate pattern generation to promote the diversity of the results. Without loss of generality, for each person, a set of $C$ ($C=10$) candidate patterns, denoted as $\mathcal{CP}$, are generated according to the historical data and $C$ candidate personas, respectively. We provide the details of these candidate personas in Appendix C.4.
@@ 3.1.2 Pattern Evaluation with Self-Consistency
This step involves assessing the consistency of the candidate patterns to identify the most plausible one. We implement a scoring mechanism to evaluate the alignment of candidate patterns with historical data. To achieve this objective, we define a scoring function to gauge each candidate pattern $cp$ in the set $\mathcal{CP}$. This function evaluates $cp$ against two distinct sets of activity trajectories: the specific activity trajectories $\mathcal{T}_{i}$ of a targeted resident $i$ and the sampled activity trajectories from other residents $\mathcal{T}_{\sim i}$:
$score_{cp}=\sum_{t\in\mathcal{T}_{i}}r_{t}-\sum_{t^{\prime}\in\mathcal{T}_{\sim i}}r_{t^{\prime}},$ (1)
where we design an evaluation prompt to ask the LLM to generate rating scores $r_{t}$ and $r_{t^{\prime}}$. Specifically, the LLM agent is prompted to assess the degree of preference for a given trajectory based on the candidate pattern. Ideally, the LLM agent should assign a higher $r_{t}$ for data from the targeted resident and a lower $r_{t^{\prime}}$ for data from other residents. This scheme essentially identifies the self-consistent pattern: the activity pattern derived from the activity trajectory data of the target user should be consistent with the data from this person during the evaluation. We provide the pseudo-code of the algorithm for Phase 1 of LLMob in Appendix A.
@@ 3.2 Motivation-Driven Activity Generation
In Phase 2 of LLMob, we focus on the retrieval of motivation and the integration of motivation and activity patterns for individual activity trajectory generation. Since the context length is limited for the LLMs, we can not expect that the LLMs can consume all the available historical information and give plausible output. Retrieval-augmented generation has been identified as a crucial factor in boosting the performance of LLM [37]. This enhancement provides additional information that aids LLM in more effectively responding to queries. While previous studies on activity generation mainly overlook the critical factors of macro temporal information (e.g., date) or specific scenarios (e.g., harsh weather) [42], we propose a more sophisticated activity generation which accounts for various conditions by taking advantage of the human-like intelligence of LLM. For instance, the activity trajectory at date $d$ can be inferred given the motivation of this date and the habitual activity pattern as:
@@
This generation scheme instructs the LLM agent to simulate a designated individual according to a given activity pattern, and then meticulously generate an activity trajectory in accordance with the daily motivation. To obtain insightful and reliable motivations toward different aspects of data availability and sufficiency, two retrieval schemes are proposed. Notably, we considered them as two promising directions for designing solutions to real-world applications, rather than claming which is superior. The detail of each retrieval scheme is introduced as follows:
@@ 3.2.1 Evolving-based Motivation Retrieval
This scheme is related to the intuitive principle that an individual’s motivation on any given day is influenced by her interests and priorities in preceding days [28]. Guided by this understanding, our approach harnesses the intelligence of the LLM agent to understand the behavior of daily activities and the underlying motivations. As illustrated in Figure 3, for a specific date $d$ for which we aim to generate the activity trajectory, we consider the activities of the past $k$ days ($k=\min(7,l)$, where $l$ is the maximum value such that the trajectory for date $d-l$ can be found in the database), and prompt the LLM agent to act as an urban resident based on the pattern identified in Section 3.1 and summarize $k$ motivations behind these activities. Using these summarized motivations, the LLM agent is further prompted to infer potential motivation for the target date $d$.
@@ 3.2.2 Learning-based Motivation Retrieval
$sim_{d_{a},d_{b}}=\sum_{t=1}^{N_{d}}\mathbf{1}_{(\mathcal{T}_{d_{a}}(t)=\mathcal{T}_{d_{b}}(t))}\text{ if }|\mathcal{T}_{d_{a}}|>t\text{ and }|\mathcal{T}_{d_{b}}|>t,$ (3)
@@
    $\mathcal{L}(\theta)=\sum_{n=1}^{N}-\log\left(\frac{e^{\text{logits}_{i}}}{\sum^{k+1}_{j=1}e^{\text{logits}_{j}}}\right)_{n},$ (5)
@@
Upon training a similarity score approximation mode, it can be applied to access the similarity between any given query date and historical dates. This enables us to retrieve the most similar historical data, which is prompted to the LLM agent to generate a summary of the motivations prevalent at that time. By doing so, we can extrapolate a motivation relevant to the query date, providing a basis for the LLM agent to generate a new activity trajectory.
@@ 4.1 Experimental Setup
Dataset. We investigate and validate LLMob over a personal activity trajectory dataset from Tokyo. This dataset was obtained through Twitter and Foursquare APIs and covers the data from January 2019 to December 2022. The time frame of this dataset is insightful as it captures typical daily life prior to the COVID-19 pandemic (i.e., normal period) and subsequent alterations during the pandemic (i.e., abnormal period). To facilitate a cost-efficient and detailed analysis for different periods, we randomly choose 100 users to model their individual activity trajectory at a 10-minute interval according to the number of available trajectories. Samples are shown in the following Table 1.
Table 1: Samples of personal activity data.
UserID Latitude Longitude Location Name Category Time
44673 35.008 139.015 Convenience Store Shop & Service 2019-12-17 8:00
44673 35.009 139.018 Ramen Restaurant Food & Service 2019-12-17 8:30
44673 35.004 139.060 Italian Restaurant Food & Service 2019-12-17 11:20
44673 35.009 139.085 Farmers Market Shop & Service 2019-12-17 14:20
44673 35.005 139.086 Soba Restaurant Food & Service 2019-12-17 18:00
We utilize the category classification in Foursquare to determine the activity category for each location. We use 10 candidate personas (Appendix C.4) as a prior for subsequent pattern generation, which captures a diverse range of activity patterns within the data of this study. For the application to other datasets, this style of candidate patterns can be easily initialized using an LLM.
[…]
After extracting the above characteristics from both the generated and real-world trajectory data, Jensen-Shannon divergence (JSD) is employed to quantify the discrepancy between them. Lower JSD is preferred.